I am trying to help my son with Equations, but I don't know how my self. Can you help?> x+2x+5=x+19
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

DON'T UNDERSTAND THE LEADING>????
IF THATIS REMOVED THEN:
X+2X+5=X+19
X+2X-X=19-5
2X=14
X=14/2
X=7 ANSWER.
PROOF
7+2*7+5=7+19
7+14+5=26
26=26

Answer by stanbon(75887)

x+2x+5=x+19
Just think of the "x" as "car"
-----------------
You have
car + 2 cars + 5 = car + 19
-----------
Substract a car from both sides to get
2cars + 5 = 19
Subtract 5 from both sides to get:
2cars = 14
Divide both sides by 2 to get:
car = 7
----------------
the same procedure would get you x = 7
===========================
Cheers,
Stan H.